OpenStudy (anonymous):

Convert 86000 cm to km. Show all of your work.

OpenStudy (shubha):

100 cm = 1 m => 1 cm = 1/100 m so, 86000 cm = 86000 * (1/100) m = 86000/100 m = 860 m now, 1000 m = 1 km => 1 m = 1/1000 km so, 860 m = 860 * (1/1000) km = 860/1000 km = 0.86 km so, 86000 cm = 0.86 km

OpenStudy (anonymous):

Hence, 86000 cm are 0.86 km. 86000 divided by 100000!

OpenStudy (anonymous):

1cm = 10 ^ -2 m 1km = 10 ^ 3 m difference is 10 ^ 5 since the unit is getting bigger the number has to be smaller -> devide by the difference 86000 / 10^5 = 0.86 (or 8.6 * 10^ -1)